OpenStudy (anonymous):

Write three different radical expressions equivalent to 3√54*a^6.

OpenStudy (mertsj):

Is it this: \[3\sqrt{54a^6}\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[3\sqrt{9\times6a^6}=3\times3\sqrt{6a^6}=9a^3\sqrt{6}\]